half the sum of two numbers is 15. half their difference is 8.what is the numbers

Answer by mananth(16949)

1/2(x+y)=15
x+y =15*2
x+y=30.......1
1/2(x-y)=8
x-y =8*2
x-y=16.......2
Add (1) & (2)
2x=46
/2
x= 23
y=7
.....